A store sells bags of dog food. Let X be the number of bags sold to the next customer. The probability distribution for X is given below. Find the variance for the number of bags sold. Give your answer to four decimal places and do not give units.

x 1 2 3 4 5
P(X=x) 0.52 0.37 0.06 0.04 0.01
Answer #1
X P(X=x) X2 X * P(X=x) X2 * P(X=x)
1 0.52 1 0.52 0.52
2 0.37 4 0.74 1.48
3 0.06 9 0.18 0.54
4 0.04 16 0.16 0.64
5 0.01 25 0.05 0.25
Total 1 55 1.65 3.43

Mean  μ = ΣΧ* P(x)) = 1.65

Variance σ = Σ(X2 * P(X) – (μ):

o2 = 3.43 - (1.65)

02 = 0.7075